solve using addition method
x - 2y = -1 and -x + 5y = 4

Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

Adding corresponding sides, we get 3y = 3, or y = 1.
Substituting into the 1st equation, x - 2 = -1, or x = 1.

Answer by mananth(16946)

x - 2y = -1
-x + 5y = 4
add up
-2y+5y=-1+4
+3y=3
/3
y=1
..
plug value of y in first equation
x-2=-1
x=+2-1
x=1
...
m.ananth@hotmail.ca